Can someone answer and explain these two problems for me please?
1. Solve this equation. Round to the nearest hundredth. e^2x-1=438
2. Fully simplify into a single logarithmic statement. log(a)b+log(a)c-log(a)d
Thank you in advance

Answer by stanbon(75887)

1. Solve this equation. Round to the nearest hundredth.
e^2x-1=438
Take the natural log of both sides to get::
2x-1 = ln(438)
2x-1 = 6.08
2x = 7.08
x = 3.54
---------------
2. Fully simplify into a single logarithmic statement.
log(a)b+log(a)c-log(a)d
----
= loga(ac/d)
===================
Cheers,
Stan H.
-------------